# bibliography.bib

### v1

@@ Abstract
In our previous notations, we have $\mathfrak{hc}=\mathfrak{hc}_{B}^{G},\chi=\chi_{B}^{G}$. It was proved in loc. cit. that if $P\subset Q\subset R$, we have
$\chi_{Q}^{R}\chi_{P}^{Q}=\chi_{P}^{R},\mathfrak{hc}^{Q}_{P}\mathfrak{hc}^{R}_{Q}=\mathfrak{hc}^{R}_{P}.$
@@
$\iota_{P}:\tilde{\Delta}_{P}\to\mathcal{Y}_{P},$
be the corresponding embedding.
Identify
@@
via the map $t\mapsto(t,tU)$. Here $P$ acts on $L$ via the adjoint action of $L$ on itself. Abusing notation, we will write $\iota_{P}^{!}$ for the composed functor
@@
where the last arrow is the forgetful functor, and let $\iota_{P*}$ similarly stand for its right adjoint.
We have the following description of the parabolic restriction $\operatorname{Res}_{P}^{G}$ functor (cf. [bezrukavnikovParabolicRestrictionPerverse2021], Remark 4.2):
@@
and parabolic induction
$\operatorname{Ind}_{P}^{G}=\chi_{P}^{G}\iota_{P*},$ (5)
We will also need the following proposition, parallel to [lusztigCharacterSheaves1985], Lemma 6.6 and [ginzburgAdmissibleModulesSymmetric1989], where it was proved for $Q=G$. For a pair of parabolics $P\subset Q$ with Levi subgroups $L$ and $M$, respectively, define
@@
In particular, identity functor is a direct summand of $\chi_{P}^{Q}\mathfrak{hc}^{Q}_{P}$.
We sketch the proof, parallel to the proof in loc. cit.
@@
To undburden the notations, we omit the quotients by the left diagonal $G$-action everywhere. All arrows in all diagrams are $G$-equivariant. By base change, left hand side is naturally isomorphic to the functor $p_{*}q^{!}$ defined by the diagram
${X}$${\mathcal{Y}_{Q}}$${\mathcal{Y}_{Q}}$$\scriptstyle{p}$$\scriptstyle{q}$
@@
$q(x_{1}U_{Q},x_{2}U_{Q},x_{3}U_{Q},x_{4}U_{Q},zP)=(x_{1}U_{Q},x_{2}U_{Q})$
$p(x_{1}U_{Q},x_{2}U_{Q},x_{3}U_{Q},x_{4}U_{Q},zP)=(x_{3}U_{Q},x_{4}U_{Q}).$
Similarly, right hand side is naturally isomorphic to the functor $p^{\prime}_{*}q^{\prime!}$ defined by the diagram
@@
$\pi(x_{1}U_{Q},x_{2}U_{Q},x_{3}U_{Q},x_{4}U_{Q},zP)=(x_{3}x_{4}^{-1}x_{2}U_{Q},x_{1}U_{Q},x_{2}U_{Q},zU_{P})\simeq\\ \simeq(x_{3}U_{Q},x_{1}x_{2}^{-1}x_{4}U_{Q},x_{4}U_{Q},zU_{P})\operatorname{mod}L.$
It is easy to see that $\pi$ is an isomorphism with the inverse $X^{\prime}\to X$
$(a,b,c,z)\mapsto(b,c,a,c,z),$
@@
We say that $\mathcal{F}\in D^{b}_{G}(\mathcal{Y}_{P})$ has character $\lambda\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$, if $\mathfrak{hc}_{B}^{P}(\mathcal{F})$ is a monodromic complex with monodromy in the $W_{J}$-orbit of ${\lambda}$. We write $D^{\lambda}(\mathcal{Y}_{P})$ for the full subcategory consisting of complexes with character $\lambda$. Note that, by definition, $D^{\lambda}(\mathcal{Y}_{P})=D^{w\lambda}(\mathcal{Y}_{P})$ for $w\in W_{J}.$ Let $P^{\lambda}(\mathcal{Y}_{P})$ stand for the full subcategory of perverse objects in $D^{\lambda}(\mathcal{Y}_{P})$. It is standard that for $X\in D^{\lambda}(\mathcal{Y}_{P}),Y\in D^{\mu}(\mathcal{Y}_{P}),\mu\notin W\lambda$ we have $\operatorname{Hom}(X,Y)=0$.
For $P=G,\mathcal{Y}_{G}\simeq G,$ objects in $P^{\lambda}(G)$ are called character sheaves with character $\lambda$.
For $\lambda\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$ let $W_{\lambda}$ be its stabilizer in $W$. Since $G$ is assumed to have connected center, $W_{\lambda}$ is a Coxeter group conjugate to a parabolic subgroup of $G$. Let $\mathfrak{m}_{\lambda+}\subset\mathfrak{m}_{0}$ be the ideal generated by $W_{\lambda}$-invariant polynomials in $\mathfrak{m}_{0}$. Let $\mathfrak{E}_{n}^{\lambda}$ be the perverse local system on $T$ satisfying $\mathcal{M}(\mathfrak{E}_{n}^{\lambda})=m_{\lambda}^{*}(\mathfrak{R}/\mathfrak{m}_{\lambda+}^{n})$. Abusing notation, write $\mathfrak{E}_{n}^{\lambda}$ for a sheaf on $\tilde{\Delta}_{B}$ corresponding to $\mathfrak{E}_{n}^{\lambda}\in D^{b}(T)$. We will need the following
@@
$\mathfrak{hc}_{*}(E_{n}^{\lambda})=\operatorname{Res}_{B}^{G}(E_{n}^{\lambda})=\oplus_{\mu\in W\lambda}\mathfrak{E}_{n}^{\mu}.$
@@
This follows from Theorem 7.1 together with Proposition 3.2 of [chenConjecturesBravermanKazhdan2019] (see also Section 5.3 from [hhh]). ∎
This allows to formulate the following analogue of Proposition 3.1.4 for categories of character sheaves on $\textrm{GL}_{n}$.
@@
Let $G=\textrm{GL}_{n}$. Let $S$ be the (finite) set of $W$-fixed points in $\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$. Write $({}^{p}D^{\leq 0},{}^{p}D^{\geq 0})$ for the perverse t-structure on $D^{b}_{G}(G)$.
@@
    For $\mathcal{F}\in D^{b}_{G}(G)$ we have $\mathcal{F}\in{{}^{p}D^{\geq 0}}$ if and only if the system ${}^{p}\mathcal{H}^{i}(\mathcal{F}\star E_{n}^{\lambda})$ is essentially zero for all $\lambda\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$ and all $i<0$.
@@
    The system ${}^{p}\mathcal{H}^{i}(\mathcal{F}\star E_{n}^{\lambda})$ is essentially zero for all $\lambda\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell}),\lambda\notin S$, and all $i$ if and only if ${}^{p}\mathcal{H}^{i}(\mathcal{F})$ are character sheaves with characters in $S$.
@@
    If for $\mathcal{F}\in P_{G}(G)$ we have ${}^{p}\mathcal{H}^{0}(\mathcal{F}\star E_{1}^{\lambda})=0$ for all $\lambda\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell}),\lambda\notin S$, then $\mathcal{F}$ is a character sheaf with characters in $S$.
@@
$\operatorname{Res}_{B}^{G}(\mathcal{F}\star E_{n}^{\lambda})=\operatorname{Res}_{B}^{G}(\mathcal{F})\star\left(\oplus_{\mu\in W\lambda}\mathcal{E}_{n}^{\mu}\right).$ (6)
By the result of [bezrukavnikovParabolicRestrictionPerverse2021], the parabolic restriction functor is t-exact as a functor
@@
Since there are no non-trivial cuspidal pairs for $D^{b}_{G}(G)$, see [lusztigIntersectionCohomologyComplexes1984],
@@
are shifted $S$-monodromic local systems. Since, for a shifted $S$-monodromic local system $\mathcal{E}$ on $T$, $\operatorname{Ind}_{B}^{G}(\mathcal{E})$ is a character sheaf with character in $S$, we get the result. ∎
@@
is an isomorphism for all $n$ and $\lambda\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell}),\lambda\notin S$, the cohomology of the cone $C_{f}$ of $f$ are character sheaves with characters in $S$.
@@
[…]
we get that ${}^{p}\mathcal{H}^{0}(C_{f})\star^{0}E_{n}^{\lambda}=0$ for $\lambda\notin S$, and so, by Proposition 3.3.3 (c), ${}^{p}\mathcal{H}^{0}(C_{f})$ is a character sheaf with character in $S$. It follows that ${}^{p}\mathcal{H}^{0}(C_{f})\star E_{n}^{\lambda}=0$ for $\lambda\notin S$, and so ${}^{p}\mathcal{H}^{-1}(C_{f}\star E_{n}^{\lambda})={}^{p}\mathcal{H}^{-1}(C_{f})\star^{0}E_{n}^{\lambda}$. Since maps $c_{\lambda}$ are surjective, we get from Proposition 3.3.3 (a) applied to $\mathcal{F}=\mathcal{B}$ that the system ${}^{p}\mathcal{H}^{-1}(C_{f})\star^{0}E_{n}^{\lambda}$ is essentially zero for all $n,\lambda\neq 0$, and so, by 3.3.3 (b), ${}^{p}\mathcal{H}^{-1}(C_{f})$ is a character sheaf with characters in $S$. ∎
@@
[…]
For objects $X,\{C_{i}\}_{i=1}^{k}$ of a triangulated category, we write $X\in\langle C_{1},C_{2},...,C_{k}\rangle$ if there exists a sequence of objects $\{X_{i}\}_{i=1}^{k},X_{1}=C_{1},X_{k}=X$, such that for all $i<k$, $(X_{i-1},X_{i},C_{i})$ is a distinguished triangle.
The following Proposition is a straightforward consequence of standard distinguished triangles for 6 functors for constructible sheaves.
@@
Let $X$ be a variety stratified by locally closed subvarieties $\{S_{t}\}_{t=1}^{n}$, and let $j_{t}:S_{t}\to X$ be the corresponding locally-closed embeddings. Assume that $S_{k}\subset\overline{S}_{l}$ implies $k<l$. Then for every $\mathcal{F}\in D^{b}(X)$, $\mathcal{F}\in\langle j_{t*}j^{!}_{t}\mathcal{F}\rangle_{t=1}^{n}$.
@@
Standard, see for example proof of Lemma 6.2.5 in [hhh]. ∎
@@
$\Pi_{\lambda}\mathfrak{hc}_{P*}(\mathcal{F})=\Pi_{\lambda}\iota_{P*}\operatorname{Res}^{G}_{P}(\mathcal{F}).$
@@
By (4) , we have $\iota_{P*}\operatorname{Res}^{G}_{P}(\mathcal{F})=\iota_{P*}\iota_{P}^{!}\mathfrak{hc}_{P*}(\mathcal{F})$. Let
@@
$(\iota_{P*}\operatorname{Res}^{G}_{L}(\mathcal{F}),\mathfrak{hc}_{P}(\mathcal{F}),j_{*}j^{*}\mathfrak{hc}_{P}(\mathcal{F})).$
@@
where $j^{\prime}$ is an embedding of an open subset
@@
where $w$ runs through minimal length representatives of non-unital congruence classes in $W/W_{J}$, and $g_{w}$ are some lifts of $w$ to $N_{G}(T)$. Let $\tilde{j}^{\prime}_{w}:\tilde{V}_{w}\to\tilde{V}$ be the corresponding locally closed embeddings. Consider $j^{\prime}_{*}\mathcal{G}$ as a $T$-equivariant complex on $G/U\times G/U$. By Proposition 3.3.5, $j_{*}^{\prime}\mathcal{G}\in\langle\tilde{j}^{\prime}_{w*}\tilde{j}_{w}^{\prime!}j_{*}^{\prime}\mathcal{G}\rangle_{w}$, so it is enough to prove that $\Pi_{\lambda}\tilde{j}^{\prime}_{w*}\tilde{j}_{w}^{\prime!}j_{*}^{\prime}\mathcal{G}=0$. Consider a point $(xU,yU)\in\tilde{V}_{w}$. Choose $x,y$ in the corresponding congruence classes so that $x^{-1}y=g_{w}l$, $l\in L.$ Since $\lambda$ is $P$-generic, there is a coweight $\mu^{\vee}:\mathbb{G}_{m}\to T$ with $\langle\lambda,\mu^{\vee}-w(\mu^{\vee})\rangle\neq 0$, and such that $\mu^{\vee}(\mathbb{G}_{m})$ is in the center of $L$. Then, for $t\in\mathbb{G}_{m}$, we have
@@
so that the $G$-equivariant sheaf on $\tilde{V}_{m}$, equivariant with respect to the right diagonal $T$-action, is also equivariant with respect to the action of $\mathbb{G}_{m}$ defined by the coweight $(\mu^{\vee}-w(\mu^{\vee}))$ on one of the components. Since non-zero equivariant complex can not have non-trivial monodromy, the result follows. ∎
@@
Assume that ${\lambda}\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$ is $P$-generic for a parabolic $P$. Restricted to the category $P^{\lambda}(G)$, the functor $\Pi_{\lambda}\operatorname{Res}_{P}^{G}$ is a braided monoidal equivalence of categories.
@@
and so
$\Pi_{\lambda}\mathfrak{hc}_{P}^{G}\chi_{P}^{G}\mathcal{F}\simeq\mathcal{F}.$
It follows that, for any $\mathcal{F}\in D^{\lambda}(L)$,
@@
is an isomorphism. This follows from [lusztigCharacterSheavesIV1986][17.12].
Since monodromic sheaves with different monodromies are orthogonal with respect to convolution, it follows from Proposition 3.3.6 that $\Pi_{\lambda}\iota_{P*}\operatorname{Res}^{G}_{P}$ is monoidal, since $\mathfrak{hc}_{P*}$ and $\Pi_{\lambda}$ are monoidal.
The functor $\mathfrak{hc}_{*}$ is equipped with the central structure, see [ginzburgAdmissibleModulesSymmetric1989], [bezrukavnikovCharacterDmodulesDrinfeld2012]. It is easy to see that, for complexes $\mathcal{F}$ with $\mathfrak{hc}_{*}(\mathcal{F})$ supported on $\tilde{\Delta}_{P}$, this central structure restricts to the braiding structure coming from $P$-equivariance.
@@
We recall some of the notations from the Introduction. Let again $P$ be an arbitrary parabolic. Define
@@
Let now $G=\textrm{GL}_{n}=\textrm{GL}(V)$, for a fixed $n$-dimensional $\mathbf{k}$-vector space $V$, let $B$ be a Borel subgroup, and let $P\supset B$ be a parabolic subgroup fixing a line in $\mathbb{P}(V)$. Write
$\mathcal{S}_{n}:=\mathcal{S}_{P}.$
We consider $\mathcal{S}_{n}$ as an object in $P_{\textrm{GL}_{n}}(\textrm{GL}_{n}).$ By Proposition 3.2.1 this is a symmetric monoidal category. Our main result is the following
@@
Exterior powers of $\mathcal{S}_{n}$ satisfy the following:
  1. a)
    $\wedge^{k}\mathcal{S}_{n}=\mathcal{S}_{\wedge^{k}},$ for $k=0,\dots,n$.
  2. b)
    $\wedge^{k}\mathcal{S}_{n}=0$ for $k>n$.
@@
$\mathcal{S}_{n}\star(-)=\chi_{P}^{G}\circ\mathfrak{hc}_{P}^{G}(-),$
@@
$\iota_{P*}\iota_{P}^{!}\mathfrak{hc}_{P}^{G}(-)\to\mathfrak{hc}_{P}^{G}(-).$
Assume that the statement of the Theorem 3.4.1 is known for $n$ and $k$. Define the morphism
@@
[…]

### v2

@@ Abstract
In our previous notations, we have $\mathfrak{hc}=\mathfrak{hc}_{P}^{G},\chi=\chi_{P}^{G}$ (note that $\mathcal{Y}_{G}\simeq G$} via the map $(x,y)\mapsto yx^{-1}$). It was proved in ibid. that, if $P\subset Q\subset R$, we have
$\chi_{Q}^{R}\chi_{P}^{Q}=\chi_{P}^{R}\text{ and }\mathfrak{hc}^{Q}_{P}\mathfrak{hc}^{R}_{Q}=\mathfrak{hc}^{R}_{P}.$
@@
$\iota_{P}:\tilde{\Delta}_{P}\to\mathcal{Y}_{P}$
be the corresponding embedding. Identify
@@
via the map $t\mapsto(t,tU)$. Here $P$ acts on $L$ through the adjoint action of $L$ on itself. Abusing notation, we will write $\iota_{P}^{!}$ for the composed functor
@@
where the last arrow is the forgetful functor. Let $\iota_{P!}$ stand for its left adjoint.
We have the following description of the parabolic restriction functor $\operatorname{Res}_{P}^{G}$ and parabolic induction functor $\operatorname{Ind}_{P}^{G}$ (cf. [BY21, Remark 4.2]):
@@
$\operatorname{Ind}_{P}^{G}$ $=\chi_{P}^{G}\iota_{P!}.$ (5)
We will also need the following proposition, parallel to [Lus04, Lemma 6.6] and, for $Q=G$, [Gin89, Lemma 8.5.4]. For a pair of parabolics $P\subset Q$ with Levi subgroups $L$ and $M$, respectively, define
@@
In particular, the identity functor is a direct summand of $\chi_{P}^{Q}\mathfrak{hc}^{Q}_{P}$.
We sketch the proof, parallel to the proof in ibid.
@@
To unburden the notations, we omit the quotients by the left diagonal $G$-action everywhere. All arrows in all diagrams are $G$-equivariant. By base change, the left-hand side is naturally isomorphic to the functor $p_{*}q^{!}$ defined by the diagram
[…]
@@
$q((x_{1}U_{Q},x_{2}U_{Q})M,(x_{3}U_{Q},x_{4}U_{Q})M,zP)$ $=(x_{1}U_{Q},x_{2}U_{Q})M,$
$p((x_{1}U_{Q},x_{2}U_{Q})M,(x_{3}U_{Q},x_{4}U_{Q})M,zP)$ $=(x_{3}U_{Q},x_{4}U_{Q})M.$
Note that the condition $x_{1}x_{2}^{-1}x_{4}x_{3}^{-1}\in{}^{z}U_{P}$ in the definition of $X$ can be rewritten as:
$z^{-1}x_{1}(x_{2}^{-1}x_{4})x_{3}^{-1}z\in U_{P}.$
Since $U_{P}\subset Q$, this and the two conditions $x_{1}^{-1}z\in Q$ and $x_{3}^{-1}z\in Q$ imply that $x_{2}^{-1}x_{4}\in Q$. We will use this observation in the computations below.
Similarly, the right-hand side is naturally isomorphic to the functor $p^{\prime}_{*}q^{\prime!}$ defined by the diagram
@@
$\pi((x_{1}U_{Q},x_{2}U_{Q})M,(x_{3}U_{Q},x_{4}U_{Q})M,zP)=\\ ((x_{3}x_{4}^{-1}x_{2}U_{Q},x_{1}U_{Q},x_{2}U_{Q})M,zP).$
It is easy to see that $\pi$ is an isomorphism with the inverse given by the following map $X^{\prime}\to X$:
$((aU_{Q},bU_{Q},cU_{Q})M,zP)\mapsto((bU_{Q},cU_{Q})M,(aU_{Q},cU_{Q})M,zP),$
@@
We say that $\mathcal{F}\in D^{b}_{G}(\mathcal{Y}_{P})$ has character $\lambda\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$, if $\mathfrak{hc}_{B}^{P}(\mathcal{F})$ is a monodromic complex with monodromy in the $W_{J}$-orbit of $\lambda$}. We write $D^{\lambda}(\mathcal{Y}_{P})$ for the full subcategory consisting of complexes with character $\lambda$. Note that, by definition, $D^{\lambda}(\mathcal{Y}_{P})=D^{w\lambda}(\mathcal{Y}_{P})$ for $w\in W_{J}.$ Let $P^{\lambda}(\mathcal{Y}_{P})$ stand for the full subcategory of perverse objects in $D^{\lambda}(\mathcal{Y}_{P})$. It is well-known that if $\mu\notin W\lambda$, then for all $X\in D^{\lambda}(\mathcal{Y}_{P})$ and $Y\in D^{\mu}(\mathcal{Y}_{P})$ we have $\operatorname{Hom}(X,Y)=0$.
For $P=G,$ where we have $\mathcal{Y}_{G}\simeq G,$ objects in $P^{\lambda}(G)$ are called character sheaves with character $\lambda$. More generally, for any finite subset $S\subset\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$, we say that objects of $\bigoplus_{\lambda\in S}P^{\lambda}(G)$, where the latter is considered as a full subcategory of $P_{G}(G),$ are character sheaves with characters in $S$.
For $\lambda\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$ let $W_{\lambda}$ be its stabilizer in $W$. Since $G$ is assumed to have connected center, $W_{\lambda}$ is a Coxeter group conjugate to a parabolic subgroup of $G$. Let $\mathfrak{m}_{\lambda+}\subset\mathfrak{m}_{0}$ be the ideal generated by $W_{\lambda}$-invariant polynomials in $\mathfrak{m}_{0}$. Let $\mathfrak{E}_{n}^{\lambda}$ be the perverse local system on $T$ satisfying $\mathcal{M}(\mathfrak{E}_{n}^{\lambda})=m_{\lambda}^{*}(\mathfrak{R}/\mathfrak{m}_{\lambda+}^{n})$. Abusing notation, write $\mathfrak{E}_{n}^{\lambda}$ for the sheaf on $\tilde{\Delta}_{B}$ corresponding to $\mathfrak{E}_{n}^{\lambda}\in D^{b}(T)$. We will need the following
@@
$\mathfrak{hc}_{*}(E_{n}^{\lambda})=\operatorname{Res}_{B}^{G}(E_{n}^{\lambda})=\bigoplus_{\mu\in W\lambda}\mathfrak{E}_{n}^{\mu}.$
@@
This follows from Theorem 7.1 together with [Che22, Proposition 3.2] (see also [BT22, Section 5.3]). ∎
This allows us to formulate the following analogue of Proposition 3.1.4 for categories of character sheaves on $\textrm{GL}_{n}$.
@@
Let $G=\mathrm{GL}_{n}$. Let $S$ be a finite $W$-invariant set of points in $\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$. Write $({}^{p}D^{\leq 0},{}^{p}D^{\geq 0})$ for the perverse $t$-structure on $D^{b}_{G}(G)$.
@@
    For $\mathcal{F}\in D^{b}_{G}(G)$ we have $\mathcal{F}\in{{}^{p}D^{\geq 0}}$ if and only if the system of $^{p}\mathcal{H}^{i}(\mathcal{F}\star E_{n}^{\lambda})$ is essentially zero for all $\lambda\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$ and all $i<0$.
@@
    The system of $^{p}\mathcal{H}^{i}(\mathcal{F}\star E_{n}^{\lambda})$ is essentially zero for all $\lambda\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$ with $\lambda\notin S$, and all $i$, if and only if the $^{p}\mathcal{H}^{i}(\mathcal{F})$ are character sheaves with characters in $S$.
@@
    If for $\mathcal{F}\in P_{G}(G)$ we have $^{p}\mathcal{H}^{0}(\mathcal{F}\star E_{1}^{\lambda})=0$ for all $\lambda\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$ with $\lambda\notin S$, then $\mathcal{F}$ is a character sheaf with characters in $S$.
@@
$\operatorname{Res}_{B}^{G}(\mathcal{F}\star E_{n}^{\lambda})=\operatorname{Res}_{B}^{G}(\mathcal{F})\star\left(\bigoplus_{\mu\in W\lambda}\mathfrak{E}_{n}^{\mu}\right).$ (6)
By the result of [BY21], the parabolic restriction functor is $t$-exact as a functor
@@
Since there are no non-trivial cuspidal pairs for $D^{b}_{G}(G)$ (see [Lus84]),
@@
are shifted $S$-monodromic local systems. For a shifted $S$-monodromic local system $\mathcal{E}$ on $T$, $\operatorname{Ind}_{B}^{G}(\mathcal{E})$ is a character sheaf with character in $S$. Since $\operatorname{Res}^{G}_{B}$ is exact and conservative, $\mathcal{F}$ is a subobject of $\operatorname{Ind}^{G}_{B}\operatorname{Res}^{G}_{B}(\mathcal{F})$, and we get the result. ∎
@@
is an isomorphism for all $n$ and $\lambda\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$ with $\lambda\notin S$, the perverse cohomology sheaves of the cone $C_{f}$ of $f$ are character sheaves with characters in $S$.
@@
$\dots\to{}^{p}\mathcal{H}^{-1}(\mathcal{A}\star E_{n}^{\lambda})\to{}^{p}\mathcal{H}^{-1}(\mathcal{B}\star E_{n}^{\lambda})\xrightarrow{c_{\lambda}}{}^{p}\mathcal{H}^{-1}(C_{f}\star E_{n}^{\lambda})\xrightarrow{0}\\ \xrightarrow{0}\mathcal{A}\star^{0}E_{n}^{\lambda}\to\mathcal{B}\star^{0}E_{n}^{\lambda}\xrightarrow{0}{}^{p}\mathcal{H}^{0}(C_{f})\star^{0}E_{n}^{\lambda}\to 0.$ (7)
We get that $^{p}\mathcal{H}^{0}(C_{f})\star^{0}E_{n}^{\lambda}=0$ for $\lambda\notin S$, and so, by Proposition 3.3.3 (c), $^{p}\mathcal{H}^{0}(C_{f})$ is a character sheaf with characters in $S$. It follows that $^{p}\mathcal{H}^{0}(C_{f})\star E_{n}^{\lambda}=0$ for $\lambda\notin S$, and so $^{p}\mathcal{H}^{-1}(C_{f}\star E_{n}^{\lambda})={}^{p}\mathcal{H}^{-1}(C_{f})\star^{0}E_{n}^{\lambda}$. Since maps $c_{\lambda}$ are surjective, we get from Proposition 3.3.3 (a) applied to $\mathcal{F}=\mathcal{B}$ that the system of $^{p}\mathcal{H}^{-1}(C_{f})\star^{0}E_{n}^{\lambda}$ is essentially zero for all $n$ whenever $\lambda\notin S$, and so, by 3.3.3 (b), $^{p}\mathcal{H}^{-1}(C_{f})$ is a character sheaf with characters in $S$. ∎
@@
We assume again that $G$ has connected center. For a complex $X\in D^{b}_{G}(\mathcal{Y}_{P})$ such that $\mathfrak{hc}_{P}(X)$ is monodromic, write $\Pi_{\lambda}X$ for the summand of $X$ with character $\lambda$. By Proposition 3.3.2, $\Pi_{\lambda}$ can be explicitly written as $\lim\limits_{\xleftarrow[n]{}}(E_{n}^{\lambda,L}\star-)$, where $E_{n}^{\lambda,L}$ denotes the sheaf constructed in Proposition 3.3.2 for $G=L$, considered as a sheaf on $\mathcal{Y}_{P}$ via the functor $\iota_{P!}$.
For objects $X,C_{1},\dots,C_{k}$ of a triangulated category, we write $X\in\langle C_{1},C_{2},...,C_{k}\rangle$ if there exists a sequence of objects $X_{1},\dots,X_{k},$ with $X_{1}=C_{1},X_{k}=X$, such that for all $i<k$ $(X_{i-1},X_{i},C_{i})$ is a distinguished triangle.
The following proposition is a straightforward consequence of standard distinguished triangles in the constructible derived categories.
@@
Let $X$ be a variety stratified by locally closed subvarieties $\{S_{t}\}_{t=1}^{n}$, and let $j_{t}:S_{t}\to X$ be the corresponding locally-closed embeddings. Assume that $S_{k}\subset\overline{S}_{l}$ implies $k<l$. Then for every $\mathcal{F}\in D^{b}(X)$ we have $\mathcal{F}\in\langle j_{t*}j^{!}_{t}\mathcal{F}\rangle_{t=1}^{n}$.
@@
See for example the proof of Lemma 6.2.5 in [BT22]. ∎
@@
$\Pi_{\lambda}\mathfrak{hc}_{P*}=\Pi_{\lambda}\iota_{P!}\operatorname{Res}^{G}_{P}.$
@@
By (4) we have $\iota_{P!}\operatorname{Res}^{G}_{P}(\mathcal{F})=\iota_{P!}\iota_{P}^{!}\mathfrak{hc}_{P*}(\mathcal{F})$. Let
@@
$(\iota_{P!}\operatorname{Res}^{G}_{L}(\mathcal{F}),\mathfrak{hc}_{P}(\mathcal{F}),j_{*}j^{*}\mathfrak{hc}_{P}(\mathcal{F})).$
@@
where $j^{\prime}$ is the embedding of an open subset
@@
where $w$ runs through minimal length representatives of non-unital congruence classes in $W/W_{J}$, and $g_{w}$ are some lifts of $w$ to $N_{G}(T)$. Let $\tilde{j}^{\prime}_{w}:\tilde{V}_{w}\to\tilde{V}$ be the corresponding locally closed embeddings. Consider $j^{\prime}_{*}\mathcal{G}$ as a $T$-equivariant complex on $G/U\times G/U$. By Proposition 3.3.5, $j_{*}^{\prime}\mathcal{G}\in\langle\tilde{j}^{\prime}_{w*}\tilde{j}_{w}^{\prime!}j_{*}^{\prime}\mathcal{G}\rangle_{w}$, so it is enough to prove that $\Pi_{\lambda}\tilde{j}^{\prime}_{w*}\tilde{j}_{w}^{\prime!}j_{*}^{\prime}\mathcal{G}=0$. Consider a point $(xU,yU)\in\tilde{V}_{w}$. Choose $x,y$ in the corresponding congruence class so that $x^{-1}y=g_{w}l$, $l\in L.$ Since $\lambda$ is $P$-generic, there is a coweight $\mu^{\vee}:\mathbb{G}_{m}\to T$ with $\langle\lambda,\mu^{\vee}-w(\mu^{\vee})\rangle\neq 0$, and such that $\mu^{\vee}(\mathbb{G}_{m})$ is in the center of $L$. Then, for $t\in\mathbb{G}_{m}$, we have
@@
and so a $G$-equivariant sheaf on $\tilde{V}_{m}$ that is equivariant with respect to the right diagonal $T$-action, is also equivariant with respect to the action of $\mathbb{G}_{m}$ defined by the coweight $\mu^{\vee}-w(\mu^{\vee})$ on one of the components. Since a non-zero equivariant complex can not have non-trivial monodromy, the result follows. ∎
@@
Assume that ${\lambda}\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$ is $P$-generic for a parabolic $P$. Restricted to the category $P^{\lambda}(G)$, the functor
$\Pi_{\lambda}\operatorname{Res}_{P}^{G}:P^{\lambda}(G)\to P^{\lambda}(L)$
is a braided monoidal equivalence of categories.
@@
and so $\Pi_{\lambda}\mathfrak{hc}_{P}^{G}\chi_{P}^{G}\mathcal{F}\simeq\mathcal{F}.$ It follows that, for any $\mathcal{F}\in D^{\lambda}(L)$,
@@
is an isomorphism. This follows from [Lus86, §17.12].
Since monodromic sheaves with different monodromies are orthogonal with respect to convolution, it follows from Proposition 3.3.6 that $\Pi_{\lambda}\iota_{P!}\operatorname{Res}^{G}_{P}$ is monoidal, since $\mathfrak{hc}_{P*}$ and $\Pi_{\lambda}$ are monoidal.
The functor $\mathfrak{hc}_{*}$ is equipped with the central structure (see [Gin89], [BFO12]). It is easy to see that, for complexes $\mathcal{F}$ with $\mathfrak{hc}_{*}(\mathcal{F})$ supported on $\tilde{\Delta}_{P}$, this central structure restricts to the braiding structure coming from $P$-equivariance, and so $\Pi_{\lambda}\operatorname{Res}_{P}^{G}$ respect the braiding. ∎
@@
We recall some of the notations from the introduction. Let $P$ again be an arbitrary parabolic. Define
@@
Let now $G=\textrm{GL}_{n}=\textrm{GL}(V)$, and $W=S_{n}$. We keep the notation of Section 1.3.2, writing $P=P_{1}$ for convenience. We consider
$\mathcal{S}_{n}:=\mathcal{S}_{P}$
as an object in $P_{\textrm{GL}_{n}}(\textrm{GL}_{n}).$ By Proposition 3.2.1, $P_{\textrm{GL}_{n}}(\textrm{GL}_{n})$ is a symmetric monoidal category. Our main result is
@@
Exterior powers of $\mathcal{S}_{n}$ with respect to the symmetric monoidal structure $\star^{0}$ satisfy the following: for $k\geq 0,$ there is an isomorphism
$\wedge^{k}\mathcal{S}_{n}\simeq\mathcal{S}_{\wedge^{k}}.$
@@
$\mathcal{S}_{n}\star(-)=\chi_{P}^{G}\mathfrak{hc}_{P}^{G}(-),$
@@
$\iota_{P!}\iota_{P}^{!}\mathfrak{hc}_{P}^{G}(-)\to\mathfrak{hc}_{P}^{G}(-).$
Assume that the statement of the Theorem 3.4.1 is known for $n$ and $k$. Define a morphism
@@
$\mathcal{S}_{\wedge^{k+1}}\xrightarrow{\alpha}\operatorname{Ind}^{G}_{P}\operatorname{Res}^{G}_{P}(\mathcal{S}_{\wedge^{k}})\xrightarrow{\tilde{\Phi}_{n}(\mathcal{S}_{\wedge^{k}})}\mathcal{S}_{n}\star^{0}\wedge^{k}\mathcal{S}_{n}\xrightarrow{\beta}\wedge^{k+1}\mathcal{S}_{n},$